OpenStudy (anonymous):

When less force is used but the amount of work done remains the same, how does distance change?

OpenStudy (btaylor):

You know that the formula for work is \(W=F \times d\). Since the amount of work is the same, let's create two formulas, one for each situation. \(W_1=F_1 \times d_1\) and \(W_2 = F_2 \times d_2\). Since you know that \(W_1\) and \(W_2\) are equal, you can set the two equations equal to each other to get \(F_1 \times d_1 = F_2 \times d_2\). From there, we can proceed with our analysis.

OpenStudy (btaylor):

We know from the problem that \(F_2<F_1\). If \(F_1 \times d_1 = F_2 \times d_2\), what does that tell us about \(d_1\) vs. \(d_2\)?

OpenStudy (agent0smith):

W = Fd If W stays the same, and F goes down, then d must go...?